OpenStudy (anonymous):

Given a cube with a volume of 36 cm3, what is the volume of a square pyramid that can fit perfectly inside the cube? 3cm3 6cm3 9cm3 12cm3

OpenStudy (anonymous):

@Gabylovesyou @bigcountry97 @bibby @midhun.madhu1987 @preetha

OpenStudy (anonymous):

yeah i know i still dont know how to get the answer

OpenStudy (midhun.madhu1987):

 volume of cube = a^3 Volume of Square Pyramid = (a^2 h)/3 the height and base of square pyramid is equal to the side of cube..